OpenStudy (anonymous):

A spinner has 10 equal sections labeled 1 through 10. A bag contains five letter tiles labeled A, A, B, B, and B. What is the probability of spinning a multiple of 3 on the spinner and drawing a letter tile labeled B from the bag?

OpenStudy (anonymous):

@DragonFlame

OpenStudy (anonymous):

Probability questions with "and" mean you need to multiply the probabilities. The probability of spinning a multiple of three is: \[\frac{3}{10}\] and the probability of drawing a 'B' is \[\frac{3}{5}\] so the probability of spining a multiple of 3 AND drawing a 'B' is \[\frac{3}{10}*\frac{3}{5}\] Which is equal to \[\frac{9}{50}\] or 18%

OpenStudy (anonymous):

thanks